OpenStudy (anonymous):

Given points A(3, 2) and B(9, 1). Answer the following question concerning using the Pythagorean Theorem to find AB. What is this value of the vertical leg?

OpenStudy (anonymous):

vertical let is 1 from 1 to 2. horizontal let is 6 from 3 oto 9. distance is \[\sqrt{1^2+6^2}=\sqrt{37}\]

OpenStudy (anonymous):

thats not an answer choice .